show that the equation of the tangent to x2+xy+y=0 at the point(x1,y1) is

(2x1+y1)x+(x1+1) y+y1 =0


1
Expert's answer
2022-02-08T12:29:07-0500
"x^2+xy+y=0"

Differentiate both sides with respect to "x"


"\dfrac{d}{dx}(x^2+xy+y)=\dfrac{d}{dx}(0)"

Use the Chain Rule


"2x+y+x\dfrac{dy}{dx}+\dfrac{dy}{dx}=0"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=-\dfrac{2x+y}{x+1}"

At the point "(x_1,y_1)"


"\dfrac{dy}{dx}|_{(x_1,y_1)}=-\dfrac{2x_1+y_1}{x_1+1}"

The equation of the tangent to "x^2+xy+y=0" at the point"(x_1,y_1)" is


"y-y_1=-\dfrac{2x_1+y_1}{x_1+1}(x-x_1)"

"(x_1+1)y+(2x_1+y_1)x-x_1y_1-y_1 -2x_1^2-x_1y_1=0"

"(x_1+1)y+(2x_1+y_1)x-2(x_1^2+x_1y_1+y_1)+y_1=0"

Then the equation of the tangent to "x^2+xy+y=0" at the point"(x_1,y_1)" will be


"(x_1+1)y+(2x_1+y_1)x+y_1=0"